Find the value of the acute angle between the lines 3x-y+1=0 and x-2y+1=0.


1
Expert's answer
2021-05-25T08:01:01-0400

Given lines are-

"3x-y+1=0\Rightarrow y=3x+1"


"x-2y+1=0\Rightarrow y=\dfrac{x}{2}+\dfrac{1}{2}"


Slope of the lines are-


"m_1=3,m_2=\dfrac{1}{2}"



So Angle between two lines-


"tan\theta=\dfrac{m_1-m_2}{1+m_2m_1}"


"=\dfrac{3-0.5}{1+3(0.5)}\\[9pt]=\dfrac{2.5}{2.5}=1"


"\theta=45^{\circ}"